A person standing on a 50 m cliff throws a rock upward at 26.5 m/s. How long will it take for the rock to reach the ground? Hint: determine the final velocity first.



1
Expert's answer
2021-07-25T09:15:59-0400

Equation of motion of the rock is given by

"y(t)=50+26.5t-\frac{9.8t^2}{2}"

At the instant when the rock reaches the ground

"y(t)=0"

Hence

"50+26.5t-\frac{9.8t^2}{2}=0"

Root

"t=6.9\:\rm s"